OpenStudy (anonymous):

OK i have a question someone has already helped me but the problem my school uses lists the answer they help me calculate as wrong can anyone else help me see if thats the softwares mistake or ours thanks the problem is 1/x^2-7x+2=1/x^2-2x+2 it states that there is a solution but we calculated it to be as undefined can someone pleas check to see if its a software issue or not thanks

OpenStudy (anonymous):

what is the question post it

OpenStudy (amistre64):

can you use better notation to clear up any ambiguity

OpenStudy (anonymous):

yeah i can post the link to the problem sorry

OpenStudy (amistre64):

this would be true of the denominators are equal, right?

OpenStudy (anonymous):

yes

OpenStudy (amistre64):

x^2-7x+2 = x^2-2x+2 -7x+2 = -2x+2 -7x = -2x -7x +2x = 0 when x=0

OpenStudy (anonymous):

ok i understand that so what would be the answer undefined or 0?

OpenStudy (amistre64):

that is when x=0 when x=0: 1/2 = 1/2 is true

OpenStudy (anonymous):

ok its 0

OpenStudy (anonymous):

It's true if and only x=0

OpenStudy (anonymous):

thanks

OpenStudy (amistre64):

youre welcome

OpenStudy (anonymous):

So, it's not undefined